# InferLog: Accelerating LLM Inference for Online Log Parsing via ICL-oriented Prefix Caching

### v1

@@ Abstract.
Modern software systems generate massive volumes of runtime logs, necessitating efficient and accurate log parsing to enable critical downstream tasks such as anomaly detection and root cause analysis. Recently, large language models (LLMs) have achieved advanced accuracy on log parsing, but their deployment in production environments faces two major limitations: (i) the privacy risks associated with commercial LLMs, driving the adoption of local deployment, and (ii) the stringent latency and throughput requirements imposed by high-volume log streams, which existing LLM-based parsers fail to meet. Although recent efforts have reduced the number of LLM queries, they overlook the high latency of the LLM invocations, where concurrent log parsing requests can cause serve performance degradation of LLM inference system.
In this study, we present InferLog, the first LLM inference optimization method for online log parsing. Our key insight is that the inference efficiency emerges as the vital bottleneck in LLM-based online log parsing, rather than parsing accuracy. InferLog accelerates inference by designing (i) A Prefix-aware ICL Refinement policy to refine the examples and permutation of in-context learning to improve the prefix caching efficiency. (ii) A rapid and task-specific configuration tuning pipeline based on meta-learning to find the optimal LLM scheduling-related configuration for dynamic log parsing workloads. The experimental results based on Loghub dataset and vLLM demonstrate that InferLog significantly outperforms existing inference optimization methods and markedly accelerates the state-of-the-art LLM-based log parser without compromising parsing accuracy.
@@ 1. Introduction
In recent years, large language models (LLMs) have revolutionized log parsing by leveraging their semantic understanding and few-shot learning capabilities(Xu et al., 2024b; Jiang et al., 2024a; Wu et al., 2025; Zhong et al., 2024; Ma et al., 2024a, b). Current LLM-based parsers achieve state-of-the-art (SOTA) accuracy in template extraction, even for unseen log formats. This advancement stands in contrast to traditional parsing methods, which typically rely on manually designed heuristics and syntactic analysis(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023; Le and Zhang, 2023; Shima, 2016). Nevertheless, a critical gap persists between existing methods and modern software system deployment: First, commercial LLMs pose significant privacy risks for log parsing due to potential unauthorized data retention and exposure vulnerabilities(Ma et al., 2024b). Consequently, organizations increasingly adopt locally deployable open-source LLMs to ensure confidential log data remains secure under strict compliance standards. Second, a modern software system can produce log data at rates of several GB per second, imposing stringent requirements on processing latency and throughput(He et al., 2021; Wang et al., 2022). Despite recent works have achieved promising parsing accuracy, the inference performance is far from the Service Level Objectives (SLOs) in production environments. In our experiments using the vLLM system, we observed that the p95 inference latency for concurrent Loghub-2k log parsing requests was in the tens of seconds, it leads to cascading delays in downstream tasks, ultimately compromising the reliability of large-scale software systems. So, it is crucial to develop an efficient local LLM inference system to handle the massive log parsing requests and meet the SLOs.
@@
Recent works(Jin et al., 2024b; Caching, 2024; Zheng et al., 2024) have highlighted the use of prefix caching (PC) technique to enhance the performance of LLM inference systems. By sharing and reusing intermediate results (i.e., KV Cache) across requests, it effectively accelerates time-to-first-token (TTFT) and reduces computation costs. While PC benefits general LLM workloads, the effectiveness in log parsing tasks remains unexplored. We conduct an empirical study of LLM-based online log parsing workloads under concurrent scenarios on vLLM(Kwon et al., 2023), three critical performance issues were identified: (i) The prefill phase (processing input sequence before generating outputs) dominates inference latency, causing 85.4% of delays distribution, violating production SLOs. (ii) Traditional prefix caching are ineffective in bypassing computations for tokens in in-context learning (ICL) examples, which make up a significant portion of the context in the prompt. (iii) LLM system parameters, such as maximum token and batch thresholds, significantly impact inference performance, through automated configuration tuning, it can lead to substantial reductions in latency compared with default setting. A detailed analysis can be found in § 2.2.
@@
Generally, There are mainly two challenges in designing the inference optimization techniques for log parsing workloads. C1: Prefix caching imposes strict requirements on same prefix tokens, as the KV cache contains unique positional embeddings within the original statement and the dependencies between the tokens of the sentence. However, log parsing prompts are dynamic across requests, characterized by varying in-context learning (ICL) examples and permutation(Xu et al., 2024b), which often leads to prefix cache misses. Existing methods(Zheng et al., 2024; Liu et al., 2024) attempt to mitigate this by reordering requests to group those with identical prefixes for batch processing, however, these approach is not suitable for online requests. C2: Rapid configuration adaptation for dynamic log parsing workloads. Logs generated by different software systems exhibit distinct characteristics and token distributions, necessitating tailored configurations for optimal performance. Existing tuning methods such as Bayesian Optimization(Cheng et al., 2024; Zhang et al., 2021; Alipourfard et al., 2017) and Reinforcement Learning(Zhang et al., 2019; Li et al., 2019; Dou et al., 2022) demand extensive online trials. This renders them impractical for online inference tasks that demand swift configuration tuning(Zhu et al., 2025).
@@
InferLog Approach. Our key insight is that the inference efficiency emerges as the vital bottleneck in LLM-based online log parsing, which is ignored in previous works(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Ma et al., 2024a; Xiao et al., 2024; Wu et al., 2025). To this end, we propose InferLog, a novel framework designed to optimize the LLM Inference performance for Log parsing tasks, focusing on reducing latency and improving throughput in online environments. InferLog proposes Prefix-aware ICL Refinement(PAIR) policy to refine the log parsing prompt by adjusting the contents and order of ICL examples. Specifically, for each current request, InferLog first identifies the examples of historical requests with the highest prefix cache hit rate probability, then performs modifying and reordering operations to update the current examples to align with it. By doing so, the ICL component can significantly boost the prefix cache hit rate(solution to C1). To achieve a fast and tailored configuration optimization, we introduce Attention mechinism(Vaswani et al., 2017) in Model-Agnostic Meta-Learning (MAML)(Finn et al., 2017) to identify optimal initial parameters of the model for the target tuning tasks and then adopt Sequential Model-based Optimization (SMBO)(Feurer et al., 2015) to update the meta-model through few-shot learning with newly collected data to rapidly recommend best configuration in a few iterations(solution to C2).
@@ 2.1.1. Log Parsing
Log parsing is the process of transforming semi-structured log entries into structured data by extracting both the static elements (i.e., log templates) and the variable components (i.e., log parameters) from the messages. Existing log parsers can be categorized into two main types: syntax-based and semantic-based. Syntax-based log parsers(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023) typically employ manually designed heuristics or analyze syntactic characteristics to extract log templates. However, their accuracy tends to decline when dealing with logs that do not adhere to established formats. In contrast, semantic-based log parsers(Liu et al., 2022; Le and Zhang, 2023; Xu et al., 2024b; Ma et al., 2024a; Jiang et al., 2024a; Xiao et al., 2024; Huang et al., 2025) utilize neural networks or language models to discern log templates and parameters by understanding the underlying meanings of log messages. For example, LogPPT(Le and Zhang, 2023) utilizes a RoBERTa model for identifying log templates and parameters through prompt-based few-shot learning. Recently, the emergence of large language models has led to the development of various LLM-based log parsers that offer enhanced log parsing capabilities and demonstrate strong adaptability to various log formats. These LLM-based parsers employ techniques such as fine-tuning(Ma et al., 2024a), prompt engineering(Jiang et al., 2024a; Xu et al., 2024b; Wu et al., 2025; Zhong et al., 2024) to optimize LLMs specifically for log parsing tasks, resulting in impressive performance.
@@ 2.1.2. LLM Inference System
LLM System Parameters. The presence of numerous configuration parameters offers flexibility for optimizing performance in inference engines(vLLM Engine Arguments, 2025). For example, in vLLM(Kwon et al., 2023), max-num-seqs and max-num-batched-tokens defines the maximum number of requests and tokens the model can process simultaneously in a single inference step. Besides, enable-prefix-caching and enable-chunked-prefill can be employed to activate respective features, which can provide significant advantages in certain scenarios. Recent study(Cheng et al., 2024) proposes optimizing parameters for vLLM engine using BO, however, this approach is inefficient for online log parsing task due to the large search space and diverse token distributions. Besides, it requires much online tuning steps, and can not easily adapt to diverse workloads. InferLog targets at the optimization of scheduling-related parameters, which play a critical role in request scheduling and resource utilization optimization in the environments with dynamic high-concurrency requests.
@@ 2.2.2. Insufficient Prefix KV Cache Reuse
Traditional prefix caching technique, as depicted in Fig.2(b)(ii), optimize LLM inference by reusing KV cache for static prompt components like common instructions. However, PC can not efficiently reuse KV cache for ICL part due to two reasons: (i) Dynamic examples selection. ICL examples are chosen based on -Nearest Neighbor (NN). For instance, LILAC(Jiang et al., 2024a) adopts Jaccard similarity and DivLog(Xu et al., 2024b) uses cosine distance to calculate the similarity between demonstration candidates and target log, leading to example sequence variations of ICL across requests. (ii) Specific permutation. Existing methods(Jiang et al., 2024a; Xu et al., 2024b) typically order the demonstrations based on their ascending similarity to the queried log, positioning the most similar demonstrations closer to it. This leads to the situation where, even if different log messages select the same examples, their order within the entire demonstration set will vary. Overall, these common practices in LLM-based log parsing impose token sequence mismatches, rendering traditional prefix caching ineffective for ICL tokens. Our experiments across 16 datasets show that for naive PC, the average prefix cache hit rate111The metric of vllm:gpu_prefix_cache_hit_rate exposed by vllm. is 55.17% when parsing distinct log messages.
However, although log messages are diverse, we identify the selection of ICL examples across different requests exhibits a notable degree of overlap, creating an opportunity for ICL to reuse historical KV cache. Specifically, (i) The candidate numbers for ICL demonstrations is limited. For instance, prior studies(Jiang et al., 2024a; Zhong et al., 2024) constrain the number of candidates to a fixed small size. (ii) The phenomenon of hotspot demonstrations. We find that specific templates from the candidate set are frequently selected during the examples selection process. As illustrated in Fig. 6, the top five hotspot demonstration templates were selected by average 36.2%, 27.2%, 24.7%, 21.2%, and 17.1% in each dataset222The x-axis ticks of Figure 5 and Figure 6 are the same as Figure 11., respectively. For instance, across 16 datasets, an average of independent log messages chose hotDemo1 as their demonstration template.
@@ 2.2.3. Suboptimal LLM Inference Parameter Configurations Setting and Inefficient Tuning Process
Configuration parameters of LLM engines significantly impacts the inference performance, particularly for scheduling-related parameters, which control the maximum batch token size and scheduling intervals, thereby determining the scheduling strategy. It is necessary to perform configuration optimization due to the following reasons. (i) Default parameters underperform for specific workloads. Default parameter settings are typically designed for general use, offering a one-size-fits-all approach that may not be optimal for specific tasks. For instance, the default max-num-batched-tokens is often set in line with the max-position-embeddings of LLM, yet this setting is usually excessive for medium or short context tasks. Our experiments in § 5.4.2 show optimized configurations can cut end-to-end latency by an average of 34.91%, and up to 57.6% compared to default settings. (ii) Impact of PC strategy on configuration governance. Under identical configurations, diverse PC strategies demonstrate distinct performance characteristics, with PAIR surpassing PC by an average of 6 seconds latency as illustrated in Fig.7(b). In particular, during the prefill phase, PAIR lowers average per-token processing time due to the improved KV cache reuse. This reduction suggests permitting more simultaneous requests or tokens to enhance GPU utilization without additional latency. Besides, the optimal configuration setting for PC are suboptimal within the context of PAIR.
Existing methods employ online learning techniques such as BO and RL to recommend configurations(Dou et al., 2022; Alipourfard et al., 2017; Zhang et al., 2019). However, these approaches are time-consuming, requiring hundreds of tuning steps. This is infeasible for online inference, where the tuning time should be minimized(Zhang et al., 2021; Sen et al., 2015). Consequently, there is a need to leverage historical knowledge to expedite the tuning process.
@@ 3.1. Overview
Specifically, the online log system generates production logs, which are processed with ICL-based prompt engineering to obtain requests. Upon requests arriving, InferLog match the historical ICL with the highest potential prefix cache hit rate from the ICL Table, then dynamically refines prompts by modifying and reordering the demonstrations of ICL to hit the prefix KV cache. To fast recommend optimal configurations, InferLog proposes an attention-based meta-learning algorithm where AttMAML trains a good initial regression model using historical tuning data, enabling rapid adaptation to new workloads. Employing AttMAML as the surrogate model, SMBO iteratively explores the parameter space, guided by the meta-learned priors to recommend near-optimal configurations within few iterations.
@@ 3.2.1. Prefix-Aware ICL Refinement
Prefix-based Matching. By maintaining the ICL Table to store historical ICL of each request(i.e., the request has completed inference and its KV cache has been stored in the system), which contains a demonstration set: , where is an example consisting of a labeled pair. In PAIR, we first traverse the ICL Table for finding the most matching DS with maximum prefix-matching count (PMC), as this is considered to achieve optimal KV cache reuse for the current DS. For example, consider the scenario of Fig.9, although both and share two examples with the same template with the . From the aspect of prefix, the first in does not appear in , so the PMC for is 0, indicating no possibility for prefix cache reuse of . In contrast, the first two in appear in current DS (i.e, IPV4 Addr: <*> and ARPT: <*>), resulting in a PMC of 2. Therefore, we mark as the for prefix sharing and align the current DS with .
Modifying. Although and share similar examples with same , they do not strictly adhere to identical prefix tokens due to variations in . Hence, it is essential to modify the current examples based on to ensure a complete cache hit. For example, the current DS includes the log: ARPT: 700311 and IPV4 Addr: 10.105.160.95, while contains log: ARPT: 699911 and IPV4 Addr: 10.105.160.98. These logs adhere to the template ARPT: <*> and IPV4 Addr: <*>. Since already has the KV cache in the memory of the LLM system, we modify the current example to align with the example from that shares the same template (indicated by the red dashed box). This operation enables examples with the same template to share the KV cache.
This operation is predicated on the assumption that demonstrations with the same template exhibit high similarity. This is because they provide semantic information that enables the LLMs to distinguish between the template and variable while remaining insensitive to specific variable values, as noted in (Min et al., 2022). This observation is further supported by the hidden representations, which yield an ultra-high cosine similarity of more than 0.99 for the representations of logs with the same template in most cases, as discussed in § 5.4.1.

### v3

@@ Abstract.
Modern software systems generate massive volumes of runtime logs, necessitating efficient and accurate log parsing to enable critical downstream tasks such as anomaly detection and root cause analysis. Recently, large language models (LLMs) have achieved advanced accuracy on log parsing, but their deployment in production environments faces two major limitations: First, the privacy risks associated with commercial LLMs, driving the adoption of local deployment. Second, online log parsing poses stringent challenges for latency and throughput. While recent methods reduce the number of LLM queries, they overlook the inherent overhead of LLM inference where concurrent log parsing requests can lead to performance degradation in the LLM inference system.
In this study, we present InferLog, the first LLM inference optimization method for online log parsing. Our key insight is that the inference efficiency emerges as the vital bottleneck in LLM-based online log parsing, rather than parsing accuracy. InferLog accelerates inference by designing (i) A prefix-aware ICL refinement strategy to refine the examples and permutation of in-context learning to improve the prefix caching efficiency. (ii) A rapid and task-specific configuration tuning pipeline based on meta-learning to find the optimal LLM inference system configuration. The experimental results based on Loghub-2k dataset and vLLM demonstrate that InferLog significantly outperforms existing inference optimization methods and markedly accelerates the state-of-the-art LLM-based log parsers without compromising parsing accuracy.
@@ 1. Introduction
In recent years, large language models (LLMs) have revolutionized log parsing by leveraging their semantic understanding and few-shot learning capabilities(Xu et al., 2024b; Jiang et al., 2024a; Wu et al., 2025; Zhong et al., 2024; Ma et al., 2024a, b). Current LLM-based parsers achieve state-of-the-art (SOTA) accuracy in template extraction, even for unseen log formats. This advancement stands in contrast to traditional parsing methods, which typically rely on manually designed heuristics and syntactic analysis(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023; Le and Zhang, 2023; Shima, 2016). Nevertheless, a critical gap persists between existing methods and modern software system deployment: First, commercial LLMs pose significant privacy risks for log parsing due to potential unauthorized data retention and exposure vulnerabilities(Ma et al., 2024b). Consequently, organizations increasingly adopt locally deployable open-source LLMs to ensure confidential log data remains secure under strict compliance standards. Second, a modern software system can produce log data at rates of several GB per second, imposing stringent requirements on processing latency and throughput(He et al., 2021; Wang et al., 2022). Despite recent works have achieved promising parsing accuracy, the inference performance is far from the Service Level Objectives (SLOs) in production environments. In our experiments using the vLLM system, we observed that the p95 inference latency for concurrent log parsing requests was in the tens of seconds, it leads to cascading delays in downstream tasks, ultimately compromising the reliability of large-scale software systems. So, it is crucial to develop an efficient local LLM inference system to handle the massive log parsing requests and meet the SLOs.
@@
Recent works(Jin et al., 2024b; Caching, 2024; Zheng et al., 2024) have highlighted the use of prefix caching (PC) to enhance the performance of LLM inference systems. By sharing and reusing intermediate results (i.e., KV cache) across requests, it effectively accelerates time-to-first-token (TTFT) and reduces computation costs. While PC benefits many LLM workloads(Zheng et al., 2024; zhu2025subgcache; pan2025kvflow), the effectiveness in log parsing tasks remains unexplored. We conduct an empirical study of LLM-based online log parsing workloads under concurrent scenarios on vLLM(Kwon et al., 2023), three critical performance issues were identified: (i) The prefill phase (processing input sequence before generating outputs) dominates inference latency, causing 85.4% of delays distribution, violating production SLOs. (ii) PC is ineffective in bypassing computations for tokens in in-context learning (ICL), which make up a significant portion of the context in the prompt. (iii) LLM system parameters, such as maximum tokens and batches, significantly impact inference performance, through automated configuration tuning, it can lead to substantial reductions in latency compared with default setting.
@@
Generally, There are mainly two challenges in designing the inference optimization techniques for log parsing workloads. C1: Prefix caching imposes strict requirements on same prefix tokens, as the KV cache contains unique positional embeddings within the original statement and the dependencies between the tokens of the sentence. However, log parsing prompts are dynamic across requests, characterized by varying in-context learning (ICL) examples and permutation(Xu et al., 2024b), which often lead to prefix cache missing. C2: Rapid configuration adaptation for dynamic log parsing workloads. Logs generated by different software systems exhibit distinct characteristics and token distributions, necessitating tailored configurations for optimal performance. Existing tuning methods such as Bayesian Optimization(Cheng et al., 2024; Zhang et al., 2021; Alipourfard et al., 2017) and Reinforcement Learning(Zhang et al., 2019; Li et al., 2019; Dou et al., 2022) demand extensive online trials. This renders them impractical for online inference tasks that demand swift configuration tuning(Zhu et al., 2025).
@@
InferLog Approach. Our key insight is that the inference efficiency emerges as the vital bottleneck in LLM-based online log parsing, which was ignored in previous works(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Ma et al., 2024a; Xiao et al., 2024; Wu et al., 2025). To this end, we propose InferLog, a novel framework designed to optimize the LLM Inference performance for Log parsing tasks, focusing on reducing latency and improving throughput in online environments. InferLog proposes Prefix-aware ICL Refinement(PAIR) policy to refine the log parsing prompt by adjusting the contents and order of ICL examples. Specifically, for each current request, InferLog first identifies the examples of historical requests with the highest prefix cache hit rate probability, then performs modifying and reordering operations to update the current examples to align with it. By doing so, the ICL component can significantly boost the prefix cache hit rate(solution to C1). To achieve a fast and tailored configuration optimization, we introduce Attention mechinism(Vaswani et al., 2017) in Model-Agnostic Meta-Learning (MAML)(Finn et al., 2017) to identify optimal initial parameters of the model for the target tuning tasks and then adopt Sequential Model-based Optimization (SMBO)(Feurer et al., 2015) to update the meta-model through few-shot learning with newly collected data to rapidly recommend high-performance configuration in a few iterations(solution to C2).
@@ 2.1.1. Log Parsing
Log parsing is the process of transforming semi-structured log entries into structured data by extracting both the static elements (i.e., log templates) and the variable components (i.e., log parameters) from the messages. Existing log parsers can be categorized into two main types: syntax-based and semantic-based. Syntax-based log parsers(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023) typically employ manually designed heuristics or analyze syntactic characteristics to extract log templates. However, their accuracy tends to decline when dealing with logs that do not adhere to established formats. In contrast, semantic-based log parsers(Liu et al., 2022; Le and Zhang, 2023; Xu et al., 2024b; Ma et al., 2024a; Jiang et al., 2024a; Xiao et al., 2024; Huang et al., 2025) utilize neural networks or language models to discern log templates and parameters by understanding the underlying meanings of log messages. For example, LogPPT(Le and Zhang, 2023) utilizes a RoBERTa model for identifying log templates and parameters through few-shot learning. Recently, the emergence of LLMs has led to the development of various LLM-based log parsers that offer enhanced log parsing capabilities and demonstrate strong adaptability to various log formats. These LLM-based parsers employ techniques such as fine-tuning(Ma et al., 2024a), prompt engineering(Jiang et al., 2024a; Xu et al., 2024b; Wu et al., 2025; Zhong et al., 2024) to optimize LLMs specifically for log parsing tasks, resulting in impressive performance.
@@ 2.1.2. LLM Inference System
LLM System Parameters. The presence of numerous configuration parameters offers flexibility for optimizing performance in inference engines(vLLM Engine Arguments, 2025). For example, in vLLM(Kwon et al., 2023), max-num-seqs and max-num-batched-tokens defines the maximum number of requests and tokens the model can process simultaneously in a single inference step. Besides, enable-prefix-caching and enable-chunked-prefill can be employed to activate respective features, which can provide significant advantages in certain scenarios. Recent study(Cheng et al., 2024) proposes optimizing parameters for vLLM engine using Bayesian Optimization (BO), however, this approach is inefficient for online log parsing task due to the large search space and diverse token distributions. Besides, it requires much online tuning steps, and can not easily adapt to diverse workloads. InferLog targets at the optimization of scheduling-related parameters, which play a critical role in requests scheduling and resources utilization optimization in the environments with dynamic and continual requests.
@@ 2.2.2. Insufficient Prefix KV Cache Reuse
Traditional prefix caching technique, as depicted in Fig.2(b)(ii), optimize LLM inference by reusing KV cache for static prompt components like common instructions. However, PC can not efficiently reuse KV cache for ICL part due to two reasons: (i) Dynamic examples selection. ICL examples are chosen based on -Nearest Neighbor (NN). For instance, LILAC(Jiang et al., 2024a) adopts Jaccard similarity and DivLog(Xu et al., 2024b) uses cosine distance to calculate the similarity between demonstration candidates and target log, leading to example sequence variations of ICL across requests. (ii) Specific permutation. Existing methods(Jiang et al., 2024a; Xu et al., 2024b) typically order the demonstrations based on their ascending similarity to the queried log, positioning the most similar demonstrations closer to it. This leads to the situation where, even different log messages select the same examples, their order within the entire demonstration set will vary. Overall, these common practices in LLM-based log parsing impose prefix token sequence mismatches, rendering traditional PC ineffective for ICL tokens. Our experiments across 16 datasets show that for naive PC, the average prefix cache hit rate111The metric of vllm:gpu_prefix_cache_hit_rate exposed by vllm. is 55.17% when parsing distinct log messages.
However, although log messages are diverse, we identify the selection of ICL examples across different requests exhibits a notable degree of overlap, creating an opportunity for ICL to reuse historical KV cache. Specifically, (i) The candidate numbers for ICL demonstrations is limited. For instance, prior studies(Jiang et al., 2024a; Zhong et al., 2024) constrain the number of candidates to a fixed small size. (ii) The phenomenon of hotspot demonstrations. We find that specific templates from the candidate set are frequently selected during the examples selection process. As illustrated in Fig. 6, the top five hotspot demonstration templates were selected by average 36.2%, 27.2%, 24.7%, 21.2%, and 17.1% in each dataset222The x-axis ticks of Figure 5 and Figure 6 are the same as Figure 11., respectively.
@@ 2.2.3. Suboptimal LLM Inference Parameter Configurations Setting and Inefficient Tuning Process
Configuration parameters of LLM engines significantly impacts the inference performance, particularly for scheduling-related parameters, which control the maximum batch token size and scheduling intervals, thereby determining the scheduling strategy. It is necessary to perform configuration optimization due to the following reasons. (i) Default parameters underperform for specific workloads. Default parameter settings are typically designed for general use, offering a one-size-fits-all approach that may not be optimal for specific tasks. For instance, the default max-num-batched-tokens is often set in line with the max-position-embeddings of LLM, yet this setting is usually excessive for medium or short context tasks. Our experiments in § 5.4.2 show optimized configurations can cut end-to-end latency by an average of 34.91%, and up to 57.6% compared to default settings. (ii) Impact of PC strategy on configuration governance. Under same configurations, diverse strategies demonstrate distinct performance characteristics, with PAIR cuts 6 s of latency versus PC in Fig.7(b). By reusing KV cache in the prefill phase, it shortens per-token time, letting us pack more requests or tokens without extra latency to enhance GPU utilization.
Existing methods employ online learning techniques such as BO and RL to search configurations(Dou et al., 2022; Alipourfard et al., 2017; Zhang et al., 2019). However, these approaches are time-consuming, requiring hundreds of tuning steps. This is infeasible for online inference, where the tuning time should be minimized(Zhang et al., 2021; Sen et al., 2015). Consequently, there is a need to leverage historical knowledge to expedite the tuning process.
@@ 3.1. Overview
Specifically, the software systems generate production logs, which are processed with ICL-based prompt engineering to obtain requests. Upon requests arriving, InferLog match the historical ICL with the highest potential prefix cache hit rate from the ICL Table, then dynamically refines prompts by modifying and reordering the demonstrations of ICL to hit the prefix KV cache. To fast recommend optimal configurations, InferLog proposes an attention-based meta-learning algorithm where AttMAML offline trains a good initial regression model using historical tuning data, enabling rapid adaptation to new workloads. Employing AttMAML as the surrogate model, SMBO iteratively explores the parameter space, guided by the meta-learned priors to recommend near-optimal configurations within few iterations.
@@ 3.2.1. Prefix-Aware ICL Refinement
Prefix-based Matching. By maintaining the ICL Table to store historical demonstration set(DS) of ICL in each request(i.e., the request has completed inference and its KV cache have been stored in the system), which contains a demonstration set: , where is an example consisting of a labeled pair. In PAIR, we first traverse the ICL Table for finding the most matching DS with maximum prefix-matching count (PMC), as this is considered to achieve optimal KV cache reuse for the current DS. For example, consider the scenario of Fig.9, although both and share two examples with the same template with the . From the aspect of prefix, the first in does not appear in , so the PMC for is 0, indicating no possibility for prefix cache reuse of . In contrast, the first two in appear in current DS (i.e, IPV4 Addr: <*> and ARPT: <*>), resulting in a PMC of 2. Therefore, we mark as for prefix sharing and align current DS with .
Modifying. Although and share similar examples with same , they do not strictly adhere to identical prefix tokens due to variations in . Hence, it is essential to modify current DS based on to ensure hitting prefix KV cache. For example, the current DS includes the log: ARPT: 700311 and IPV4 Addr: 10.105.160.95, while contains log: ARPT: 699911 and IPV4 Addr: 10.105.160.98. These logs adhere to the template ARPT: <*> and IPV4 Addr: <*>. Since already has the KV cache in the memory of the LLM system, we modify the current example to align with the example from that shares the same template (indicated by the red dashed box). This operation enables examples with the same template to share the KV cache.
This operation is predicated on the assumption that demonstrations with the same template exhibit high similarity. This is because they provide semantic information that enables the LLMs to distinguish between the template and variable while remaining insensitive to specific variable values, as noted in (Min et al., 2022). This observation is further supported by the hidden representations, which yield a high cosine similarity for the representations of logs with the same template in most cases, as discussed in § 5.4.1.